OpenStudy (anonymous):

Whis of the following ordered pairs below is not on the same line as the other three ordered pairs? Why? (2,-3), (-2,2), (0,-1), (4,-7)

OpenStudy (anonymous):

compute some slopes. see which match

OpenStudy (anonymous):

first two slope is \[\frac{2-(-3)}{-2-2}=-\frac{5}{4}\]

OpenStudy (anonymous):

second two \[\frac{2-(-1)}{0-(-2)}=\frac{3}{2}\]\]

OpenStudy (anonymous):

first and third \[\frac{-1-(-3)}{0-2}-1\]\]

OpenStudy (anonymous):

first and 4th also -1 \[\frac{-7+3}{4-2}=-1\]\]

OpenStudy (anonymous):

meant -2 sorry

OpenStudy (anonymous):

first and third, first and fourth have slope -2. point (-2,2) not on the line

OpenStudy (anonymous):

all others have slope -2